OpenStudy (anonymous):

Mr. Jenkins can dig a ditch in one hour less then his son. Working together, they can do the job in 2\3 hour. How long does it take for Mr. Jenkins to do the job alone?

OpenStudy (turingtest):

J=time for Jenkins to dig a ditch S= time for his son to dig a ditch 1/(J-1)=1/S 1/(J+S)=2/3 sub in for S from the first equation: 2/3=1/(J+S)=1/(2J-1) 2J-1=3/2 2J=5/2 J=5/4 hrs

OpenStudy (turingtest):

or 1 hr 15 min